# Text-Free Multi-domain Graph Pre-training: Toward Graph Foundation Models

### v3

@@ Abstract
Given the ubiquity of graph data, it is intriguing to ask: Is it possible to train a graph foundation model on a broad range of graph data across diverse domains? A major hurdle toward this goal lies in the fact that graphs from different domains often exhibit profoundly divergent characteristics. Although there have been some initial efforts in integrating multi-domain graphs for pre-training, they primarily rely on textual descriptions to align the graphs, limiting their application to text-attributed graphs. Moreover, different source domains may conflict or interfere with each other, and their relevance to the target domain can vary significantly. To address these issues, we propose MDGPT, a text free Multi-Domain Graph Pre-Training and adaptation framework designed to exploit multi-domain knowledge for graph learning. First, we propose a set of domain tokens to to align features across source domains for synergistic pre-training. Second, we propose a dual prompts, consisting of a unifying prompt and a mixing prompt, to further adapt the target domain with unified multi-domain knowledge and a tailored mixture of domain-specific knowledge. Finally, we conduct extensive experiments involving six public datasets to evaluate and analyze MDGPT, which outperforms prior art by up to 37.9%.
1 Introduction
@@
Graph data are ubiquitous in many domains, owing to their ability to model complex relationships among entities. Thus, applications of graph data are widespread, ranging from e-commerce graphs for recommendation systems [1, 2], to citation graphs for bibliographic study [3, 4], and social networks for user behavioral analysis [5, 6]. These graphs constitute a vast knowledge repository with abundant and comprehensive information across various domains. Toward building a foundation model for graph analysis [7], it is vital to train a universal graph model based on a broad range of graph data from multiple domains, which can be further adapted to solving diverse graph-centric tasks across different downstream fields.
@@
Figure 1: Motivation of MDGPT. (a) Different scenarios of pre-training and downstream adaptation. (b) Accuracy of one-shot node classification with DGI on two target domains, namely, Cora and Photo, as more source domains are added to pre-training.
@@
However, mainstream graph learning methods fall short of achieving this goal. Conventional approaches often train graph neural networks (GNNs) [8, 9, 10] or graph transformers [11, 12, 13] in a supervised fashion, which requires not only re-training for each specific task, but also significant labeled data for each task. Inspired by the advances in pre-trained models for language [14, 15] and vision [16, 17] data, pre-training graph models [18, 19, 20] has been widely explored to overcome the limitations of supervised approaches. They learn task-irrelevant, universal properties on unlabeled graphs, and the pre-trained graph model can be further tailored to different downstream tasks using some task-specific labels through finetuning [19, 20] To further narrow the task gap between pre-training and downstream tasks, prompt learning methods [21, 22] emerge as a popular alternation of finetuning approaches. However, in prevailing solutions to graph pre-training the pre-training and downstream graphs typically originate from the same dataset consisting of one or more graphs in the same domain, including different subgraphs of a large graph [19, 21], or a collection of similar graphs [18, 23].
@@
Therefore, it is imperative to pre-train graph models on a broad spectrum of multi-domain graphs. Yet, graphs from different domains may exhibit profoundly divergent characteristics. For example, the small-world properties of a social network are not applicable to molecular graphs. The distinct nature of graphs from various domains poses a genuine hurdle in integrating broad multi-domain knowledge and achieving universal cross-domain transfer. On one hand, although some works have attempted cross-domain transfer from a single source domain [24, 25, 26, 27, 28], they fail to leverage comprehensive knowledge from a broad range of domains. On the other hand, a few recent studies [29, 30] employ large language models to process or extract node features based on associated textual descriptions in multi-domain graphs, where texts act as a unifying medium across different domains. Consequently, these approaches are limited to text-attributed graphs [31], and do not extend to general graphs that lack explicit textual data. In this work, we propose MDGPT, a foundational graph framework for Text-Free Multi-Domain Graph Pre-Training, to facilitate downstream adaptation to both seen and unseen domains. The solution is non-trivial due to two key challenges.
@@
First, how do we align multi-domain graphs in the pre-training phase? Node features in various domains exhibit different distributions, often with unique dimensionality and semantics. For example, features in a citation network denote words in a paper, whereas in a commerce graph, they describe attributes of a product, as shown in Fig. 1(a). Therefore, directly integrating multi-domain graphs may result in conflicts or interferences rather than synergy. As illustrated in Fig. 1(b), adding more source domains often leads to worse performance for conventional graph pre-training methods such as DGI [19]. Although some techniques like singular value decomposition (SVD) [32] can be used to align feature dimensions, the semantic meanings remain misaligned. A recent study [33] proposes a virtual node for each domain, linking all nodes within and then interlinking these virtual nodes. However, this method lacks an explicit mechanism for semantic alignment, leading to suboptimal performance. In MDGPT, we propose a series of domain tokens to align node features from different domains. Each domain-specific token is associated with a source domain, serving to bridge the semantic meanings in distinct domains. The token takes the form of a learnable vector, encouraging a synergistic multi-domain integration.
@@
Second, how do we adapt multi-domain prior knowledge to downstream tasks in different domains? The downstream task could be in either a seen or unseen domain. A recent multi-domain pre-training work [33] employs a finetuning or prompt-based approach for downstream adaptation. However, multi-domain prior knowledge contains not only unified global knowledge across domains, but also domain-specific knowledge in each domain. Conventional adaptation approaches for single-domain graph data fall short of effectively transferring such prior knowledge to the target domain. In MDGPT, we propose dual prompts, consisting of a unifying prompt and a mixing prompt, to align the downstream target domain with the source domains, thus facilitating the transfer of multi-domain knowledge to downstream tasks. On one hand, unifying prompt is a learnable vector, serving as a bridge to holistically align the target domain with the unified pre-trained knowledge from all source domains. On the other hand, mixing prompt integrates the individual domain tokens, aligning the target domain with a tailored mixture of knowledge from each source domain, thereby obtaining domain-specific prior knowledge for more precise adaptation.
@@
2 Related Work
Pre-training methods on graphs [18, 19, 21, 34, 35] capitalize on the intrinsic properties of graph structures via self-supervised learning. Pre-trained knowledge can be transferred to downstream tasks through fine-tuning or parameter-efficient adaptation. However, these methods assume that the pre-training and downstream graphs come from the same domain (or very similar domains), such as different subgraphs of a large graph [20, 35], or a collection of similar graphs in the same field [18, 23]. As a result, these methods struggle to generalize across multi-domain graphs.
Another line of research on cross-domain learning addresses multiple domains [24, 25, 26, 27]. These studies focus on transferring prior knowledge pre-trained on a single source domain to a different target domain by exploiting domain-invariant representations. However, they are only pre-trained on a single source domain, falling short of utilizing comprehensive multi-domain knowledge. Additionally, they are often tailored to specific tasks or domains [24, 25, 26, 27], or requiring domain expertise [24, 26], which prevents them from generalize to diverse domains and tasks.
Toward multi-domain graph pre-training, OFA [29] and HiGPT [30] employ a language model to derive and process node features from different domains through the medium of natural language, thereby limiting themselves to text-attributed graphs [34, 36]. For text-free graphs, GCOPE [33] introduces a series of domain-specific, interconnecting virtual nodes, which link to other nodes from each domain. While these virtual nodes facilitate some level of connection and alignment across the multiple domains, they do not explicitly align node features from different domains, and thus cannot effectively overcome domain conflicts. Moreover, it only transfers unified pre-trained knowledge to the target domain, overlooking source domain-specific knowledge. Inspired by multi-task pre-training [37, 38], in MDGPT, we propose the notions of domain tokens and dual prompts to overcome the shortcomings of existing multi-domain graph pre-training methods. It is important to note that multi-task pre-training addresses the interference arising from various pre-training tasks within a single domain, which is distinct from our goal to mitigate domain conflicts involving multiple domains.
3 Preliminaries
@@
Multi-domain pre-training. Each pre-training graph belongs to a source domain $D_{S_{i}}$, characterized by a node distribution $P^{(V)}_{S_{i}}$ and edge distribution $P^{(E)}_{S_{i}}$. Hence, given a set of multi-domain pre-training graphs $\mathcal{G}_{S}$ and domains $\mathcal{D}_{S}$, we have $\{(G_{1},D_{S_{1}}),(G_{2},D_{S_{2}}),\ldots,(G_{K},D_{S_{K}})\}$, where each graph $G_{i}\in\mathcal{G}_{S}$ belongs to domain $D_{S_{i}}\in\mathcal{D}_{S}$. Note that more than one graph may belong to the same domain, i.e., we may have $S_{i}=S_{j}$ for some $i,j\leq K$. In general, different domains follow different node and edge distributions, with varying node feature dimensions and semantics.
@@ 4 Proposed Approach
Figure 2: Overall framework of MDGPT.
4.1 Overall framework
@@
Pre-training. As shown in Fig. 2(a), we first align the feature dimensions of graphs from multiple source domains. Next, we propose a series of domain tokens to further unify the semantic spaces of features from different domains. Subsequently, we pre-train a graph encoder using a self-supervised task based on a universal task template following previous work [21].
@@
4.2 Multi-domain pre-training
@@
$\tilde{\vec{X}}_{i}=\mathtt{DA}_{S_{i}}(\vec{X}_{i}),$ (1)
where $\mathtt{DA}_{S_{i}}\colon\mathbb{R}^{|V|\times d_{S_{i}}}\rightarrow\mathbb{R}^{|V|\times\tilde{d}}$ is the dimension alignment function specific to domain $D_{S_{i}}$. Here $d_{S_{i}}$ is the feature dimension in domain $D_{S_{i}}$, and $\tilde{d}$ is the aligned dimension that is uniform across all domains. Various choices for the alignment function exit, such as singular value decomposition (SVD) [32], or a learnable multi-layer perceptron (MLP) [39]. In our work, each alignment function $\mathtt{DA}_{S_{i}}$ is implemented using SVD. However, dimension alignment only transform the feature spaces into the same dimensions, yet they still occupy distinct semantic spaces across different domains. Direct pre-training on such dimension-aligned feature matrices from multiple domains would result in interference rather than synergy.
@@
$\vec{H}_{S}=\mathtt{GraphEncoder}(\mathcal{G}_{S},\mathcal{X}_{S};\Theta),$ (3)
where each row of $\vec{H}_{S}$ corresponds to the output embedding of a node found in the collection of source domain graphs $\mathcal{G}_{S}$. More precisely, let $\vec{h}_{S,v}$ denote the embedding vector of node $v$, which is a node in one of the source domain graphs.
Pre-training loss. Following previous work [21, 40], we employ link prediction as the pre-training task, leveraging the abundant links in multi-domain graphs as self-supervision. In particular, we adopt the universal task template based on subgraph similarity [21], which enables compatibility among common graph-based tasks, including link prediction, node classification and graph classification. The compatibility of the pre-training task would facilitate adaptation to different downstream tasks.
Specifically, the pre-training data consist of triplets taking the form $(v,a,B)$, where $v,a$ are nodes from the source domain graphs in $\mathcal{G}_{S}$, and $B=\{b_{1},\ldots,b_{m}\}$ represents a set of nodes from $\mathcal{G}_{S}$. Here, $(v,a)$ is an edge present in $\mathcal{G}_{S}$, while $B$ serves as the negative examples such that $(v,b_{i})$ is not an edge, $\forall b_{i}\in B$. Such triplets can be readily sampled from $\mathcal{G}_{S}$ to generate the pre-training dataset $\Omega_{\text{pre}}$. Subsequently, utilizing the universal task template [21], we define the pre-training loss as
$\textstyle\mathcal{L}_{\text{pre}}(\Omega_{\text{pre}};\Theta,\mathcal{T})=-\sum_{(v,a,B)\in\Omega_{\text{pre}}}\ln\frac{\exp\left(\frac{1}{\tau}\mathtt{sim}(\mathbf{h}_{S,v},\mathbf{h}_{S,a})\right)}{\sum_{b_{i}\in B}\exp\left(\frac{1}{\tau}\mathtt{sim}(\mathbf{h}_{S,v},\mathbf{h}_{S,b_{i}})\right)},$ (4)
@@
4.3 Downstream adaptation
In addition to addressing the task gap between pre-training and downstream tasks [21, 40], we focus on transferring pre-trained multi-domain knowledge to downstream tasks, aiming to overcome the domain gap. Likewise, the downstream target domain, $D_{T}$, may have different feature dimensions from the aligned dimensions of the source domains. Therefore, the first step is to apply the same dimension alignment method as used in the pre-training phase. Given a downstream graph $G=(V,E,\vec{X})\in\mathcal{G}_{T}$ from target domain $D_{T}$, we transform its feature matrix into $\tilde{\vec{X}}=\mathtt{DA}_{T}(\vec{X})$.

### v4

@@ Abstract.
Given the ubiquity of graph data, it is intriguing to ask: Is it possible to train a graph foundation model on a broad range of graph data across diverse domains? A major hurdle toward this goal lies in the fact that graphs from different domains often exhibit profoundly divergent characteristics. Although there have been some initial efforts in integrating multi-domain graphs for pre-training, they primarily rely on textual descriptions to align the graphs, limiting their application to text-attributed graphs. Moreover, different source domains may conflict or interfere with each other, and their relevance to the target domain can vary significantly. To address these issues, we propose MDGPT, a text free Multi-Domain Graph Pre-Training and adaptation framework designed to exploit multi-domain knowledge for graph learning. First, we propose a set of domain tokens to to align features across source domains for synergistic pre-training. Second, we propose a dual prompts, consisting of a unifying prompt and a mixing prompt, to further adapt the target domain with unified multi-domain knowledge and a tailored mixture of domain-specific knowledge. Finally, we conduct extensive experiments involving six public datasets to evaluate and analyze MDGPT, which outperforms prior art by up to 37.9%.(Codes are available at https://anonymous.4open.science/r/MDGPT for anonoymous review.)
Data mining, graph learning, graph foundation model, multi-domain pre-training, prompt learning, few-shot learning.
Corresponding authors.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Web miningccs: Information systems Data miningccs: Computing methodologies Learning latent representations
1. Introduction
@@
Graph data are ubiquitous in many domains, owing to their ability to model complex relationships among entities. Thus, applications of graph data are widespread, ranging from e-commerce graphs for recommendation systems (He et al., 2020; Wu et al., 2022), to citation graphs for bibliographic study (Kanakia et al., 2019; Xiong et al., 2017), and social networks for user behavioral analysis (Fan et al., 2019; Perozzi et al., 2014). These graphs constitute a vast knowledge repository with abundant and comprehensive information across various domains. Toward building a foundation model for graph analysis (Liu et al., 2023a), it is vital to train a universal graph model based on a broad range of graph data from multiple domains, which can be further adapted to solve diverse graph-centric tasks across different downstream fields.
@@
Figure 1. Motivation of MDGPT. (a) Different scenarios of pre-training and downstream adaptation. (b) Accuracy of one-shot node classification with DGI on two target domains, namely, Cora and Photo, as more source domains are added to pre-training.
@@
However, mainstream graph learning methods fall short of achieving this goal. Conventional approaches often train graph neural networks (GNNs) (Kipf and Welling, 2017; Veličković et al., 2018a; Yu et al., 2023a) or graph transformers (Hu et al., 2020a; Yun et al., 2019; Wu et al., 2024) in a supervised fashion, which requires not only re-training for each specific task, but also significant labeled data for each task. Inspired by the advances in pre-trained models for language (Dong et al., 2019; Kenton and Toutanova, 2019) and vision (Chen et al., 2021; Hangbo et al., 2022) data, pre-training graph models (Hu et al., 2020b; Veličković et al., 2018b; You et al., 2020) has been widely explored to overcome the limitations of supervised approaches. They learn task-irrelevant, universal properties on unlabeled graphs, and the pre-trained graph model can be further tailored to different downstream tasks using some task-specific labels through finetuning (Veličković et al., 2018b; You et al., 2020) To further narrow the task gap between pre-training and downstream tasks, prompt learning methods (Liu et al., 2023b; Sun et al., 2023) emerge as a popular alternation of finetuning approaches. However, in prevailing solutions to graph pre-training, the pre-training and downstream graphs typically originate from the same dataset consisting of one or more graphs in the same domain, including different subgraphs of a large graph (Veličković et al., 2018b; Liu et al., 2023b), or a collection of similar graphs (Hu et al., 2020b; Qiu et al., 2020).
@@
Therefore, it is imperative to pre-train graph models on a broad spectrum of multi-domain graphs. Yet, graphs from different domains may exhibit profoundly divergent characteristics. For example, the small-world properties of a social network are not applicable to molecular graphs. The distinct nature of graphs from various domains poses a genuine hurdle in integrating broad multi-domain knowledge and achieving universal cross-domain transfer. On one hand, although some works have attempted cross-domain transfer from a single source domain (Ding et al., 2021; Hassani, 2022; Wang et al., 2021, 2023; Yang and Yuen, 2019), they fail to leverage comprehensive knowledge from a broad range of domains. On the other hand, a few recent studies (Liu et al., 2024; Tang et al., 2024) employ large language models to process or extract node features based on associated textual descriptions in multi-domain graphs, where texts act as a unifying medium across different domains. Consequently, these approaches are limited to text-attributed graphs (Wen and Fang, 2023a), and do not extend to general graphs that lack explicit textual data. In this work, we propose MDGPT, a foundational graph framework for Text-Free Multi-Domain Graph Pre-Training, to facilitate downstream adaptation to both seen and unseen domains. The solution is non-trivial due to two key challenges.
@@
First, how do we align multi-domain graphs in the pre-training phase? Node features in various domains exhibit different distributions, often with unique dimensionality and semantics. For example, features in a citation network denote words in a paper, whereas in a commerce graph, they describe attributes of a product, as shown in Fig. 1(a). Therefore, directly integrating multi-domain graphs may result in conflicts or interferences rather than synergy. As illustrated in Fig. 1(b), adding more source domains often leads to worse performance for conventional graph pre-training methods such as DGI (Veličković et al., 2018b). Although some techniques like singular value decomposition (SVD) (Stewart, 1993) can be used to align feature dimensions, the semantic meanings remain misaligned. A recent study (Zhao et al., 2024) proposes a virtual node for each domain, linking all nodes within and then interlinking these virtual nodes. However, this method lacks an explicit mechanism for semantic alignment, leading to suboptimal performance. In MDGPT, we propose a series of domain tokens to align node features from different domains. Each domain-specific token is associated with a source domain, serving to bridge the semantic meanings in distinct domains. The token takes the form of a learnable vector, encouraging a synergistic multi-domain integration.
@@
Second, how do we adapt multi-domain prior knowledge to downstream tasks in different domains? The downstream task could be in either a seen or unseen domain. A recent multi-domain pre-training work (Zhao et al., 2024) employs a finetuning or prompt-based approach for downstream adaptation. However, multi-domain prior knowledge contains not only unified global knowledge across domains, but also domain-specific knowledge in each domain. Conventional adaptation approaches for single-domain graph data fall short of effectively transferring such prior knowledge to the target domain. In MDGPT, we propose dual prompts, consisting of a unifying prompt and a mixing prompt, to align the downstream target domain with the source domains, thus facilitating the transfer of multi-domain knowledge to downstream tasks. On one hand, unifying prompt is a learnable vector, serving as a bridge to holistically align the target domain with the unified pre-trained knowledge from all source domains. On the other hand, mixing prompt integrates the individual domain tokens, aligning the target domain with a tailored mixture of knowledge from each source domain, thereby obtaining domain-specific prior knowledge for more precise adaptation.
@@
2. Related Work
Pre-training methods on graphs (Hu et al., 2020b; Veličković et al., 2018b; Liu et al., 2023b; Yu et al., 2024a, 2023b) capitalize on the intrinsic properties of graph structures via self-supervised learning. Pre-trained knowledge can be transferred to downstream tasks through fine-tuning or parameter-efficient adaptation. However, these methods assume that the pre-training and downstream graphs come from the same domain (or very similar domains), such as different subgraphs of a large graph (You et al., 2020; Yu et al., 2023b), or a collection of similar graphs in the same field (Hu et al., 2020b; Qiu et al., 2020). As a result, these methods struggle to generalize across multi-domain graphs.
Another line of research on cross-domain learning addresses multiple domains (Ding et al., 2021; Hassani, 2022; Wang et al., 2021, 2023). These studies focus on transferring prior knowledge pre-trained on a single source domain to a different target domain by exploiting domain-invariant representations. However, they are only pre-trained on a single source domain, falling short of utilizing comprehensive multi-domain knowledge. Additionally, they are often tailored to specific tasks or domains (Ding et al., 2021; Hassani, 2022; Wang et al., 2021, 2023), or requiring domain expertise (Ding et al., 2021; Wang et al., 2021), which prevents them from generalize to diverse domains and tasks.
Toward multi-domain graph pre-training, OFA (Liu et al., 2024) and HiGPT (Tang et al., 2024) employ a language model to derive and process node features from different domains through the medium of natural language, thereby limiting themselves to text-attributed graphs (Yu et al., 2024a; Wen and Fang, 2023b). For text-free graphs, GCOPE (Zhao et al., 2024) introduces a series of domain-specific, interconnecting virtual nodes, which link to other nodes from each domain. While these virtual nodes facilitate some level of connection and alignment across the multiple domains, they do not explicitly align node features from different domains, and thus cannot effectively overcome domain conflicts. Moreover, it only transfers unified pre-trained knowledge to the target domain, overlooking source domain-specific knowledge. Inspired by multi-task pre-training (Wang et al., 2022; Yu et al., 2024c), in MDGPT, we propose the notions of domain tokens and dual prompts to overcome the shortcomings of existing multi-domain graph pre-training methods. It is important to note that multi-task pre-training addresses the interference arising from various pre-training tasks within a single domain, which is distinct from our goal to mitigate domain conflicts involving multiple domains.
3. Preliminaries
@@
Multi-domain pre-training. Each pre-training graph belongs to a source domain $D_{S_{i}}$, characterized by a node distribution $P^{(V)}_{S_{i}}$ and edge distribution $P^{(E)}_{S_{i}}$. Hence, given a set of multi-domain pre-training graphs $\mathcal{G}_{S}$ and domains $\mathcal{D}_{S}$, we have $\{(G_{1},D_{S_{1}}),(G_{2},D_{S_{2}}),\\ \ldots,(G_{K},D_{S_{K}})\}$, where each graph $G_{i}\in\mathcal{G}_{S}$ belongs to domain $D_{S_{i}}\in\mathcal{D}_{S}$. Note that more than one graph may belong to the same domain, i.e., we may have $S_{i}=S_{j}$ for some $i,j\leq K$. In general, different domains follow different node and edge distributions, with varying node feature dimensions and semantics.
@@ 4. Proposed Approach
Figure 2. Overall framework of MDGPT.
4.1. Overall framework
@@
Pre-training. As shown in Fig. 2(a), we first align the feature dimensions of graphs from multiple source domains. Next, we propose a series of domain tokens to further unify the semantic spaces of features from different domains. Subsequently, we pre-train a graph encoder using a self-supervised task based on a universal task template following previous work (Liu et al., 2023b).
@@
4.2. Multi-domain pre-training
@@
(1) $\tilde{\vec{X}}_{i}=\mathtt{DA}_{S_{i}}(\vec{X}_{i}),$
where $\mathtt{DA}_{S_{i}}\colon\mathbb{R}^{|V|\times d_{S_{i}}}\rightarrow\mathbb{R}^{|V|\times\tilde{d}}$ is the dimension alignment function specific to domain $D_{S_{i}}$. Here $d_{S_{i}}$ is the feature dimension in domain $D_{S_{i}}$, and $\tilde{d}$ is the aligned dimension that is uniform across all domains. Various choices for the alignment function exit, such as singular value decomposition (SVD) (Stewart, 1993), or a learnable multi-layer perceptron (MLP) (Kruse et al., 2022). In our work, each alignment function $\mathtt{DA}_{S_{i}}$ is implemented using SVD. However, dimension alignment only transform the feature spaces into the same dimensions, yet they still occupy distinct semantic spaces across different domains. Direct pre-training on such dimension-aligned feature matrices from multiple domains would result in interference rather than synergy.
@@
(3) $\vec{H}_{S}=\mathtt{GE}(\mathcal{G}_{S},\mathcal{X}_{S};\Theta),$
where $\mathtt{GE}$ represents graph encoder, and each row of $\vec{H}_{S}$ corresponds to the output embedding of a node found in the collection of source domain graphs $\mathcal{G}_{S}$. More precisely, let $\vec{h}_{S,v}$ denote the embedding vector of node $v$, which is a node in one of the source domain graphs.
Pre-training loss. Following previous work (Liu et al., 2023b; Sun et al., 2022), we employ link prediction as the pre-training task, leveraging the abundant links in multi-domain graphs as self-supervision. In particular, we adopt the universal task template based on subgraph similarity (Liu et al., 2023b), which enables compatibility among common graph-based tasks, including link prediction, node classification and graph classification. The compatibility of the pre-training task would facilitate adaptation to different downstream tasks.
Specifically, the pre-training data consist of triplets taking the form $(v,a,B)$, where $v,a$ are nodes from the source domain graphs in $\mathcal{G}_{S}$, and $B=\{b_{1},\ldots,b_{m}\}$ represents a set of nodes from $\mathcal{G}_{S}$. Here, $(v,a)$ is an edge present in $\mathcal{G}_{S}$, while $B$ serves as the negative examples such that $(v,b_{i})$ is not an edge, $\forall b_{i}\in B$. Such triplets can be readily sampled from $\mathcal{G}_{S}$ to generate the pre-training dataset $\Omega_{\text{pre}}$. Subsequently, utilizing the universal task template (Liu et al., 2023b), we define the pre-training loss as
(4) $\textstyle\mathcal{L}_{\text{pre}}(\Omega_{\text{pre}};\Theta,\mathcal{T})=-\sum_{(v,a,B)\in\Omega_{\text{pre}}}\ln\frac{\exp\left(\frac{1}{\tau}\mathtt{sim}(\mathbf{h}_{S,v},\mathbf{h}_{S,a})\right)}{\sum_{b_{i}\in B}\exp\left(\frac{1}{\tau}\mathtt{sim}(\mathbf{h}_{S,v},\mathbf{h}_{S,b_{i}})\right)},$
@@
4.3. Downstream adaptation
In addition to addressing the task gap between pre-training and downstream tasks (Liu et al., 2023b; Sun et al., 2022), we focus on transferring pre-trained multi-domain knowledge to downstream tasks, aiming to overcome the domain gap. Likewise, the downstream target domain, $D_{T}$, may have different feature dimensions from the aligned dimensions of the source domains. Therefore, the first step is to apply the same dimension alignment method as used in the pre-training phase. Given a downstream graph $G=(V,E,\vec{X})\in\mathcal{G}_{T}$ from target domain $D_{T}$, we transform its feature matrix into $\tilde{\vec{X}}=\mathtt{DA}_{T}(\vec{X})$.